OpenStudy (cherrilyn):

evaluate integral:

OpenStudy (cherrilyn):

\[\int\limits_{?}^{?}dx/x(x ^{2}+25)\]

OpenStudy (dumbcow):

use partial fractions to split into two integrals \[\frac{A}{x} + \frac{Bx+C}{x ^{2}+25} = \frac{1}{x(x ^{2}+25)}\] (A+B)x^2 +Cx+25A =1 A+B =0 -->A=-B C=0 25A=1 -->A=1/25 B=-1/25 \[=\frac{1}{25}\int\limits_{}^{}\frac{1}{x}dx - \frac{1}{25}\int\limits_{}^{}\frac{x}{x ^{2}+25} dx\] Fro second integral use u substitution: u = x^2+25 du = 2x dx -->dx=du/2x \[=\frac{1}{25}\ln x - \frac{1}{50}\int\limits_{}^{}\frac{1}{u}du\] \[=\frac{1}{25}\ln x - \frac{1}{50}\ln |x ^{2}+25| +C\]